OpenStudy (anonymous):

what is the square by square root of x2-4x=0

OpenStudy (anonymous):

\[x ^{2}-4x = 0\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(x+2\sqrt{x})(x-2\sqrt{x})\]

OpenStudy (anonymous):

I'm guessing this is what your looking for?

OpenStudy (anonymous):

well yes but in a different method like x2-32=0 x=addtion and subtraction4 root of2

OpenStudy (anonymous):

you lost me there

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\(x^2-4x=0\iff x(x-4)=0\implies x=0\) or \(x=4\)

OpenStudy (anonymous):

ah ok that makes sense. He wants to find the roots

OpenStudy (anonymous):

yes that too but, :( its not the one im looking for'

OpenStudy (anonymous):

\[x^{2}-144=0\} \[x^{2}=144\rightarrow \sqrt{x}=\pm \sqrt{144}\rightarrow \alpha =\pm12\]